OpenStudy (anonymous):

How does one solve y=x^(1/x) using Logarithmic Differentiation?

ganeshie8 (ganeshie8):

are you trying to find derivative of y ?

OpenStudy (anonymous):

d/dx of the entire equation?

ganeshie8 (ganeshie8):

then you want to find \(\large \dfrac{dy}{dx}\)

OpenStudy (anonymous):

What's the difference between d/dx vs. dy/dx?

ganeshie8 (ganeshie8):

take natural log both sides, then differentiate both sides

OpenStudy (anonymous):

I don't know how to :(

ganeshie8 (ganeshie8):

d/dx is a operator dy/dx is the derivative of y

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Can you point me to a good free resource that'll teach me how to differentiate logarithmic functions?

OpenStudy (anonymous):

ok thank you

OpenStudy (akashdeepdeb):

\[y = x^{\frac{1}{x}}\] \[\log~ y = \log~ (x^{\frac{1}{x}}) = \frac{1}{x} \log~x ~~~\because \log ~a^b = b~\log~a\] \[\frac{d}{dx} [ \log~y] = \frac{d}{dx} [\frac{1}{x} \log~x]\] Now differentiate. :)

OpenStudy (anonymous):

Could you show me how to differentiate?

myininaya (myininaya):

\[\frac{d}{dx}(\ln(x))=\frac{1}{x} \\ \frac{d}{dx}\ln(f(x))=\frac{1}{f(x)} \cdot f'(x)\]

OpenStudy (perl):